OpenStudy (aravindg):

solve the foll equation x+y=(2pi)/3, sin x/sin y =2

OpenStudy (anonymous):

\[x + y = \frac{2\pi}{3}\]\[\frac{\sin x}{\sin y}=2\] \[x = \frac{2\pi}{3} - y\] \[\frac{\sin(\frac{2\pi}{3}-y)}{\sin y} = 2\]\[\sin(\frac{2\pi}{3})\cos y-\cos(\frac{2\pi}{3})\sin y = 2 \sin y\]\[\frac{\sqrt{3}}{2}\cos y - (-\frac{1}{2})\sin y = 2\sin y\]\[\sqrt{3}\cos y + \sin y = 4 \sin y\]\[\sqrt{3}\cos y = 3\sin y\]\[\frac{\sqrt{3}}{3}=\tan y\]\[y=\tan^{-1}(\frac{\sqrt{3}}{3})\]\[y=\frac \pi 6\] \[x=\frac{2\pi}{3}-\frac \pi 6\]\[x=\frac \pi 2\]

OpenStudy (jamesj):

Nicely done Jyqft